OpenStudy (anonymous):

simplify the sum. 7/a+8 + 7/a^2-64

OpenStudy (anonymous):

\[\frac{ 7 }{ a+8 } + \frac{ 7 }{ a^{2}-64 }\]

jimthompson5910 (jim_thompson5910):

First factor the second denominator \[\Large \frac{7}{a+8} + \frac{7}{a^2-64}\] \[\Large \frac{7}{a+8} + \frac{7}{(a-8)(a+8)}\] What's the next step?

OpenStudy (anonymous):

im not sure. can you walk me through it?

jimthompson5910 (jim_thompson5910):

if you wanted to add fractions like 1/2 and 2/3, how would you do it?

OpenStudy (anonymous):

oh i know. (a-8)(a+8) would be a^2-64

jimthompson5910 (jim_thompson5910):

that is correct

jimthompson5910 (jim_thompson5910):

do you see why I factored?

OpenStudy (anonymous):

so now its\[\frac{ 7 }{ a^{2}-64 }\]

OpenStudy (anonymous):

yes i see. whats the next step,?

jimthompson5910 (jim_thompson5910):

how do you add fractions? what is the key thing needed?

OpenStudy (anonymous):

umm not sure

jimthompson5910 (jim_thompson5910):

something to do with the denominators

OpenStudy (anonymous):

idk :/

jimthompson5910 (jim_thompson5910):

have a look at this page http://www.mathsisfun.com/fractions_addition.html

jimthompson5910 (jim_thompson5910):

hopefully that's all review

OpenStudy (anonymous):

so ive gotta make the denominators equal?

jimthompson5910 (jim_thompson5910):

exactly

jimthompson5910 (jim_thompson5910):

the denominators at this point are (a+8) and (a-8)(a+8) what's missing from the first denominator?

OpenStudy (anonymous):

i though the denominators where (a+8) and (a^2-64) now...

jimthompson5910 (jim_thompson5910):

that's what they were originally

jimthompson5910 (jim_thompson5910):

but I factored a^2-64

OpenStudy (anonymous):

oh. right. sorry i was confused

jimthompson5910 (jim_thompson5910):

the denominators at this point are (a+8) and (a-8)(a+8) what's missing from the first denominator?

OpenStudy (anonymous):

(a-8) ?

jimthompson5910 (jim_thompson5910):

so you have to multiply top and bottom of the first fraction by (a-8) to get that first denominator equal to the second

jimthompson5910 (jim_thompson5910):

\[\Large \frac{7}{a+8} + \frac{7}{a^2-64}\] \[\Large \frac{7}{a+8} + \frac{7}{(a-8)(a+8)}\] \[\Large \frac{7(a-8)}{(a-8)(a+8)} + \frac{7}{(a-8)(a+8)}\]

jimthompson5910 (jim_thompson5910):

What's next?

OpenStudy (anonymous):

hmmm. does anything cancel out?

OpenStudy (anonymous):

i honestly have no idea. i cant figure it out

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

are the denominators the same at this point?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so you can add up the numerators and place that over the common denominator \[\Large \frac{7(a-8)}{(a-8)(a+8)} + \frac{7}{(a-8)(a+8)}\] \[\Large \frac{7(a-8)+7}{(a-8)(a+8)}\] I'll let you simplify from here

OpenStudy (anonymous):

\[\frac{ 7a-49 }{ (a-8)(a+8) }\]

jimthompson5910 (jim_thompson5910):

you nailed it

jimthompson5910 (jim_thompson5910):

optionally you can expand out the denominator to get \[\Large \frac{7a-49}{(a-8)(a+8)}\] \[\Large \frac{7a-49}{a^2 - 64}\]